OpenStudy (anonymous):

x^2+100X-1000 find the real solution if any. using quadratic formula.

OpenStudy (anonymous):

easier to complete the square which is the same only without the huge computation

OpenStudy (anonymous):

true, but i have to solve the equation in the form givin i assume. but not sure. could be a trick question lol.

jhonyy9 (jhonyy9):

x_1,_2=(-100 +/- sqrt(10000 +4000))/2

OpenStudy (anonymous):

\[x^2+100x-1000=0\] \[x^2+100x=1000\] \[(x-50)^2=1000+50^2=3500\] \[x+50=\pm\sqrt{3500}=\pm10\sqrt{35}\] \[x=-50\pm10\sqrt{35}\]